# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ 4.2 Forecasting accuracy
Forecasting accuracy for all tasks, models, and horizons is enumerated in Table 1. Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage of Error (sMAPE666sMAPE (Appendix D.1) can provide balanced insights when metrics are averaged across different tasks. or $\epsilon\leavevmode\nobreak\ \%$) are used as our evaluation metrics. Note that for the ILI rate prediction tasks, the error metrics represent the average MAE and sMAPE across $4$ consecutive test seasons (per season results are available in Tables D1D2, and D3, including linear correlation with the target variable as an additional metric), whereas for the remaining benchmark tasks performance metrics are obtained from one train / test split for each forecasting horizon, following previously reported evaluations.
DeformTime displays an overall superior accuracy. It outperforms the most competitive method (that might be a different one per task and horizon), reducing its MAE by $10\%$ and its sMAPE by $7.4\%$ on average across all tasks and forecasting horizons. In more absolute terms, DeformTime is the best performing model in all but one occasion (where it is the second best) based on MAE, and $21$ out of $24$ times based on sMAPE. With the exception of the weather task that draws upon a $1$-year data set from a single location and hence is only expected to offer limited insights while attempting to predict $720$ time steps ($7.5$ days) ahead (see also Appendix A.5), our method offers $>\!10\%$ of MAE reduction in the most challenging forecasting tasks. Specifically, MAE is reduced by $10.5\%$ in the ETT tasks (for $H\!=\!720$ hours), and by $14.9\%$ (England) or $10.3\%$ (US Regions) in the ILI tasks (for $H\!=\!28$ days). Therefore, performance gains do not decrease as the forecasting horizon increases.
When comparing DeformTime to a specific forecaster, we notice that MAE reduction is $>$$15\%$ on average across all tasks and forecasting horizons, ranging from a $15.2\%$ reduction vs. Crossformer to $34.8\%$ vs. DLinear. [^7] A brief note to further explain DLinear’s performance is provided in Appendix D.6. These significant performance gains highlight our model’s capacity to consistently perform well under different tasks and horizons compared to other SOTA models. Interestingly, DLinear fails to surpass the average accuracy of a naïve persistence model (DeformTime reduces persistence’s MAE by $30.1\%$).
Turning our focus to the more interpretable task of predicting ILI rates, we first notice that forecasters that model inter-variable dependencies (Crossformer, LightTS, TimeMixer, iTransformer, and DeformTime) showcase superior performance. Compared to DeformTime’s average sMAPE (across all locations) ranging from $18.96\%$ to $26.14\%$ for $H\!=\!7$ to 2828\!2828 days ahead forecasting respectively, the only consistent competitor is Crossformer ($18.87\%$ to $28.83\%$). The rest of the methods do not perform well as the task becomes more challenging, reaching average sMAPEs ranging from $35.51\%$ to $45.84\%$ for $H\!=\!28$ days. Pending a more detailed evaluation (out-of-scope for this work), there is at least partial evidence to support that DeformTime is a SOTA forecaster for ILI [36, 30, 33, 31]. It not only demonstrates a great regression fit, but also captures the overall trend while forecasting $28$ days ahead (average correlation is >.$>\!.90$, see Tables D1D2, and D3).
We depict the ILI rate forecasts of all models in Appendix G. A snapshot is presented in Figure 2, showcasing $28$ days ahead predictions for influenza season 2018/19 in England. We observe that many models resemble the naïve persistence model, providing smooth but shifted and hence uninformative forecasts. DLinear provides the overall worst fit. With the exception of the iTransformer (see Appendix B for a brief justification), competitive baseline models that capture inter-variable dependencies do perform better. However, their estimates are noisy, and either not capable of capturing the onset (TimeMixer) or intensity (Crossformer and less so LightTS) of the influenza season. Contrary to that, DeformTime provides smoother and more accurate forecasts that corroborate our choice to account for both inter- and intra-variable dependencies.
@@
Further experiments, presented in Appendix D.3, provide evidence that DeformTime is also robust to random seed initialisation (Table D4).
4.3 Ablation study for DeformTime
We perform an ablation study to understand the contribution of various operations in DeformTime, namely the V-DAB and T-DAB modules, the position embeddings used in DABs (we conventionally denote both by $\mathbf{P}_{\texttt{v},\texttt{t}}$), and the NAE module with and without position embedding $\mathbf{P}_{\texttt{n}}$. Experiments are conducted on the ILI rate forecasting task for $2$ locations, England (ILI-ENG) and US Region 9 (ILI-US9), and across the $4$ test seasons. Hyperparameters are re-tuned using the same validation approach separately for each ablation variant.
Table 2 enumerates the ablation outcomes, showing the average MAE across all test periods. Evidently, each component or operation contributes to the reduction of MAE. On average, the V-DAB module provides stronger performance improvements ($8.4\%$) compared to the T-DAB module ($4.8\%$). However, for the longest forecasting horizon ($H\!=\!28$) both modules have a similar level of contribution (respectively, $11.4\%$ and $9.8\%$). Hence, establishing inter-variable dependencies is always useful, but intra-variable dependencies become more important in longer forecasting horizons.
@@ 4.3 Ablation study for DeformTime
Table 2: Ablation study for DeformTime (shortened to DTime). We report the average MAE over the $4$ test periods for tasks ILI-ENG and ILI-US9 across all forecasting horizons ($H$). ‘$\neg$’ denotes the absence of a module or an operation.
$H$ DTime $\neg$ V-DAB $\neg$ T-DAB $\neg$ $\mathbf{P}_{\texttt{v},\texttt{t}}$$\neg$ NAE $\neg$ $\mathbf{P}_{\texttt{n}}$
@@
$H$ DTime $\neg$ V-DAB $\neg$ T-DAB $\neg$ $\mathbf{P}_{\texttt{v},\texttt{t}}$$\neg$ NAE $\neg$ $\mathbf{P}_{\texttt{n}}$
@@
The inclusion of the NAE component is equally important as it improves MAE by $6.3\%$ on average. Interestingly, the level of relative improvement increases as the forecasting horizon extends (from $5.8\%$ to $7.6\%$). This potentially highlights that longer forecasting horizons require more abstraction over the feature space, favouring learning from correlated groups of variables. Finally, the use of position embeddings, albeit not entailing a very sophisticated operation, significantly enhances the impact of the V-DAB/T-DAB and NAE modules by $7.2\%$ and $8.9\%$, respectively. We argue that position embeddings, obtained before ($\mathbf{P}_{\texttt{n}}$) and after ($\mathbf{P}_{\texttt{v},\texttt{t}}$) deformation, work in tandem to maintain key information that evidently further improves the overall forecasting accuracy.
@@ 4.4 Computational complexity and efficiency of DeformTime
The computational complexity of DeformTime is $\mathcal{O}(L^{2}d+Ld^{2})$, where $L$ denotes the look-back window and $d$ is the size of the hidden layers throughout the method (see Appendix D.5 for a complete derivation). We note that the number of operations for an encoder layer reduces quadratically as we increase $r$ (T-DAB’s time window) to values greater than $1$. Likewise, segmenting the $L$ input time steps to patches of a smaller length (in V-DAB) results in quadratic computational benefits.
Overall, DeformTime has a relatively small (GPU VRAM) memory footprint (see Figure D2). Compared to other transformer-based MTS forecasting methods (Crossformer, PatchTST, and iTransformer), our method exhibits consistently low memory consumption as the number of input variables increases (vs. an exponential increase for the other methods). Furthermore, increasing the look-back window does not overly affect memory consumption either (vs. a more accelerated linear increase for Crossformer and PatchTST). Hence, we deduce that DeformTime is adept to MTS forecasting tasks with a considerable amount of exogenous predictors and longer look-back windows.

### v3

@@ 4.2 Forecasting accuracy
Forecasting accuracy for all tasks, models, and horizons is enumerated in Table 1. Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage of Error (sMAPE777sMAPE (Appendix E.1) can provide more balanced insights when metrics are averaged across different tasks. or $\epsilon\leavevmode\nobreak\ \%$) are used as our evaluation metrics. Note that for the ILI rate prediction tasks, the error metrics represent the average MAE and sMAPE across $4$ consecutive test seasons (per season results are available in Tables S1S2, and S3, where we also include the linear correlation with the target variable as an additional performance metric), whereas for the remaining benchmark tasks performance metrics are obtained from one train / test split for each forecasting horizon, following previously reported evaluations.
@@
DeformTime displays an overall superior accuracy. It outperforms the most competitive method (that might be a different one per task and horizon), reducing MAE by $7.2\%$ and sMAPE by $4.5\%$ on average across all tasks and forecasting horizons. In more absolute terms, DeformTime is the best performing model in all but two occasions based on MAE, and $20$ out of $24$ times based on sMAPE. If we exclude the weather task that uses a $1$-year data set from a single location and hence is only expected to offer limited insights while attempting to predict $720$ time steps ($7.5$ days) ahead (see also Appendix A.4), our method offers on average $9.3\%$ of MAE reduction in the most challenging forecasting tasks (maximum forecasting horizon $H$). Specifically, MAE is reduced by $5.5\%$ in the ETT tasks (for $H\!=\!720$ hours), and by $14.9\%$ (England) or $4.4\%$ (US Regions) in the ILI tasks (for $H\!=\!28$ days). Therefore, performance gains compared to other forecasting methods do not decrease as the forecasting horizon increases.
When comparing DeformTime to a specific forecasting method (as opposed to the best performing one per task and horizon), we notice that MAE reduction is $>$$11\%$ on average across all tasks and forecasting horizons, ranging from a $11.9\%$ reduction vs. ModernTCN to $34.8\%$ vs. DLinear. [^8] A brief note to further explain DLinear’s performance is provided in Appendix E.7. These significant performance gains highlight our model’s capacity to consistently perform well under different tasks and horizons compared to other SOTA models. Interestingly, DLinear fails to surpass the average accuracy of a naïve persistence model (DeformTime reduces persistence’s MAE by $30.1\%$).
Turning our focus to the more interpretable task of predicting ILI rates, we first notice that forecasting models that do not model inter-variable dependencies (DLinear, PatchTST, and CycleNet) or do so after embedding the input along the sequence dimension (iTransformer, TimeXer) showcase an overall inferior performance. This highlights the potential benefit of incorporating variable dependencies over time steps or slices that preserve the underlying temporal patterns. Such structure is a common design choice amongst the better-performing models (DeformTime, ModernTCN, Crossformer, LightTS, and TimeMixer). Compared to DeformTime’s average sMAPE (across all locations) ranging from $18.96\%$ to $26.14\%$ for $H\!=7$ to 2828\!2828 days ahead forecasting respectively, the two consistent competitors are Crossformer ($18.87\%$ to $28.83\%$) and ModernTCN ($19.67\%$ to $29.57\%$). The rest of the methods do not perform well as the task becomes more challenging, reaching average sMAPEs ranging from $35.51\%$ (LightTS) to $50.39\%$ (TimeXer) for $H\!=\!28$ days. Pending a more detailed evaluation (out-of-scope for this work), there is at least partial evidence to support that DeformTime is a SOTA forecaster for ILI (Reich et al., 2019; Morris et al., 2021; Osthus & Moran, 2021; Morris et al., 2023). It not only demonstrates a great regression fit, but also captures the overall trend while forecasting $28$ days ahead across 3 different geographical regions in 2 countries (average correlation is >.$>\!.90$, see Tables S1S2, and S3).
We depict the ILI rate forecasts of all models in Appendix G. A snapshot is presented in Figure 2, showcasing $28$ days ahead predictions for influenza season 2018/19 in England. We observe that many models resemble the naïve persistence model that predicts the last seen value of the target variable in the autoregressive input, providing smooth but uninformative and increasingly inaccurate (as $H$ increases) forecasts. DLinear provides the overall worst fit. With the exception of TimeXer and iTransformer (see also Appendix C), competitive baseline models that capture inter-variable dependencies provide more informative predictions. However, their estimates —at least in this example— appear to be more noisy, and are either not capable of capturing the onset of the influenza season (TimeMixer) or its intensity (ModernTCN, Crossformer and less so LightTS). Contrary to that, DeformTime provides smoother and more accurate forecasts that corroborate our choice to account for both inter- and intra-variable dependencies.
Further experiments, presented in Appendix E.6, provide evidence that DeformTime is also robust to random seed initialisation (Table S5). In Appendix E.8 and Table S6, we also provide results where the forecast error is averaged over the entire output sequence as opposed to only considering predictions at the target forecasting horizon time step, with additional discussion offering some insights as to why this may not be the most appropriate evaluation approach. Nevertheless, DeformTime remains the best-performing forecasting model.
@@
Table 2: Ablation study for DeformTime. We report the average MAE over the $4$ test periods for ILI tasks across all forecasting horizons, $H=\!\{7,14,21,28\}$ days ahead. ‘$\neg$’ denotes the absence of a module or an operation. $\mathbf{P}_{\texttt{v},\texttt{t}}$ denotes both position embeddings used in DABs (V-DAB, T-DAB). $\mathbf{P}_{\texttt{n}}$ denotes the position embedding used in NAE.
$H$ DeformTime $\neg$ V-DAB $\neg$ T-DAB $\neg$ $\mathbf{P}_{\texttt{v},\texttt{t}}$$\neg$ NAE $\neg$ $\mathbf{P}_{\texttt{n}}$
@@
ILI-US2 7 0.4122 0.4600 0.4165 0.4507 0.4414 0.4612
14 0.4752 0.4796 0.4758 0.5099 0.5033 0.4973
21 0.5425 0.5790 0.5563 0.5696 0.5474 0.5545
28 0.5538 0.5924 0.5794 0.5902 0.5960 0.5881
Avg. 0.4959 0.5278 0.5070 0.5301 0.5220 0.5253
@@
4.3 Ablation analysis
We perform an ablation study to understand the contribution of various operations in DeformTime, namely the V-DAB and T-DAB modules, the position embeddings used in DABs (we conventionally denote both by $\mathbf{P}_{\texttt{v},\texttt{t}}$), and the NAE module with and without position embedding $\mathbf{P}_{\texttt{n}}$. Experiments are conducted on the ILI rate forecasting task for $2$ locations, England (ILI-ENG) and US Region 9 (ILI-US9), across all forecasting horizons ($H=\!\{7,14,21,28\}$ days ahead) the $4$ test seasons. Hyperparameters are re-tuned using the same validation approach separately for each ablation variant. Further details are provided in Appendix E.2.
Table 2 enumerates the ablation outcomes, showing the average MAE across all test periods. Evidently, each component or operation contributes to the reduction of MAE. On average, the V-DAB module provides stronger performance improvements ($7.7\%$) compared to the T-DAB module ($3.9\%$). However, for the longest forecasting horizon ($H\!=\!28$), the difference between the level of contribution from these modules is lower (respectively, $9.8\%$ and $8.0\%$). Hence, establishing inter-variable dependencies is always useful, but intra-variable dependencies appear to become more important in longer forecasting horizons.
@@ 4.3 Ablation analysis
The inclusion of the NAE component is equally important as it improves MAE by $6.4\%$ on average. Interestingly, the level of relative improvement increases as the forecasting horizon extends (from $6.7\%$ to $7.1\%$). This potentially highlights that longer forecasting horizons require more abstraction over the feature space, favouring learning from correlated groups of variables. Finally, the use of position embeddings, albeit not entailing a very sophisticated operation, significantly enhances the impact of the V-DAB/T-DAB and NAE modules by $6.4\%$ and $7.9\%$, respectively. We argue that position embeddings, obtained before ($\mathbf{P}_{\texttt{n}}$) and after ($\mathbf{P}_{\texttt{v},\texttt{t}}$) deformation, work in tandem to maintain key information that evidently further improves the overall forecasting accuracy.
Interestingly, comparing the results in Tables 2 and 1, DeformTime without the V-DAB module, which results in a simpler model that captures inter-variable dependencies via the NAE module and a GRU decoder, yields similar accuracy as ModernTCN. We consider that the added value of V-DAB stems from its ability to flexibly capture variable dependencies that may not be temporally aligned. These dependencies, if they exist, have been shown to potentially contribute to better forecasting performance when aligned (Zhao & Shen, 2024). ModernTCN, however, does not specifically model such unaligned temporal dependencies. Our analysis suggests that this omission is an important factor for the observed performance degradation.
Table 3: Forecasting accuracy for models that capture inter-variable dependencies on the ILI-ENG task under increasing amount of input (exogenous) variables. The best results are in bold font and the second best are underlined.
Models iTransformer TimeXer TimeMixer LightTS Crossformer ModernTCN DeformTime
$H$ $C$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
7 $32$ 2.5141 28.67 2.7315 32.62 2.1333 24.68 1.9263 33.12 1.6659 27.51 1.9468 37.37 1.6790 23.56
$64$ 2.3658 27.79 2.9042 33.52 2.1768 25.03 2.2359 41.94 1.9090 30.72 1.7984 41.54 1.4904 23.23
$128$ 2.3346 27.15 2.8816 33.29 2.4315 27.96 2.1155 46.60 1.6160 24.80 1.8501 31.70 1.4468 22.68
$256$ 2.3325 27.06 2.7711 33.60 2.4842 28.34 2.5989 35.87 1.8018 22.17 1.9449 33.33 1.5797 25.03
$512$ 2.4466 27.98 2.9864 35.19 2.8672 32.94 2.7775 51.29 1.7930 22.53 2.1583 46.45 1.6314 23.65
14 $32$ 3.4327 39.69 3.4781 40.86 3.1351 35.30 2.8046 38.21 2.5710 34.28 2.6098 41.93 2.1651 30.25
$64$ 3.2482 37.76 3.5437 41.27 3.2472 37.42 2.9960 48.12 2.5851 35.37 2.4487 34.03 2.0055 29.36
$128$ 3.4923 39.94 3.5598 41.56 3.3225 37.71 2.9178 40.72 2.5510 32.87 2.4755 37.70 2.0362 29.55
$256$ 3.3671 38.98 3.4928 41.95 3.3600 39.06 2.6980 47.30 3.1166 38.46 2.7400 38.99 2.1626 31.91
$512$ 3.3427 38.53 3.6936 44.14 3.7326 41.78 2.7095 39.44 3.1111 39.70 2.9224 43.48 2.2848 33.15
21 $32$ 4.6290 55.13 4.4526 54.44 4.0293 56.15 3.2743 45.51 2.8971 33.86 2.7548 41.25 2.7363 37.05
$64$ 4.1034 48.80 4.3821 51.99 3.9984 54.02 3.4519 52.31 3.0058 36.95 2.7769 51.61 2.5821 37.15
$128$ 4.2907 50.79 4.3504 51.15 3.7035 53.27 4.5949 60.69 2.8090 36.49 2.9763 40.86 2.6646 37.65
$256$ 4.4976 53.14 4.5287 56.65 3.7519 52.49 3.5740 52.22 3.0846 38.88 2.8275 41.56 2.5746 34.64
$512$ 4.2563 49.04 4.6257 56.27 3.5404 45.61 3.3258 54.37 3.6267 44.76 2.9788 49.13 2.6973 35.55
28 $32$ 4.8528 58.29 4.9145 61.07 4.8393 60.72 3.7597 46.94 3.5441 44.01 3.1705 41.17 2.8000 34.75
$64$ 4.6811 56.08 4.8851 61.35 4.5416 58.54 3.6672 57.08 3.6433 46.64 3.3598 49.85 2.7637 34.17
$128$ 4.8265 56.90 4.8329 60.65 4.7736 57.94 3.7655 56.16 3.4786 44.29 3.0837 47.69 2.7945 36.90
$256$ 5.0621 60.88 4.9689 62.48 4.4152 57.56 3.6055 53.02 3.6987 45.14 3.1944 46.01 2.9963 41.02
$512$ 5.0417 59.91 5.3472 68.10 4.0968 54.52 3.5609 46.86 4.0750 48.02 3.2858 47.23 3.1113 38.94
4.4 Assessing the impact of an increasing number of exogenous predictors
We also assess the accuracy of models as the number of exogenous predictors increases. For this we use the ILI rate forecasting task given its relatively large amount of exogenous predictors. We set the number of exogenous input variables, i.e. the frequency time series for different search queries, to $C$ to $\{32,64,128,256,512\}$. [^9] The total number of input variables is $C\!+\!1$ if we also factor in the autoregressive part of the input. Queries are selected by obtaining the top-$C$ most correlated ones with the target variable (ILI rate) in the 5 last flu seasons in the training data (a detailed data set description is provided in Appendix A.2). We hypothesise that models that consistently give better performance as $C$ increases are more robust to handling multi-variable input and capture inter-variable dependencies better. To this end, we focus our analysis on models that have components that aim to capture information across variables, namely iTransformer, TimeXer, TimeMixer, LightTS, Crossformer, ModernTCN, and DeformTime. Experiments are conducted on the ILI forecasting task for England across all $4$ test seasons and we report the average MAE and sMAPE scores. All hyperparameters, except the number of input variables (see Appendix D.2), are re-tuned for each model (for all train/test splits).
Forecasting accuracy results are enumerated in Table 3. First, considering the entirety of reported outcomes, we note that DeformTime provides superior performance, reducing the average MAE and sMAPE by $11.8\%$ and $10.6\%$ respectively compared to the best-performing baseline(s). When we consider outcomes for a consistent setting of the number of input variables, DeformTime yields the lowest MAE in all but one forecasting task (vs. Crossformer for $C\!=\!32$ and $H\!=\!7$ days ahead). This re-affirms our previously reported results and also highlights that DeformTime tends to outperform competitive baselines irrespectively of the number of exogenous predictors, maintaining its superiority as the forecasting horizon increases.
Consistent with the results presented in section 4.2, the most competitive baselines are ModernTCN and Crossformer, where DeformTime reduces their MAE by $13.9\%$ and $17.1\%$ on average, respectively. iTransformer and TimeXer are the two least competitive models, with DeformTime reducing the average MAE by $37.9\%$ and $41.9\%$ respectively. This provides some evidence for the importance of structural design when it comes to time series modelling as highlighted in section 4. Specifically, disrupting the input’s temporal patterns prior to modelling is likely to cause model collapse, resulting in learning a simplistic forecasting choice resembling the persistence model, as also reflected in Figure 2. It also underscores the need for conducting evaluation on diverse datasets and metrics to comprehensively assess model performance (TimeXer is the best-performing baseline for ETTh1 but less competitive for other tasks).
@@ 4.5 Computational complexity and efficiency of DeformTime
The computational complexity of DeformTime is $\mathcal{O}(L^{2}d+Ld^{2})$, where $L$ and $d$ respectively denote the size of the look-back window and the number of hidden layers throughout the method. A detailed derivation of this is provided in Appendix E.5. We note that the number of operations for an encoder layer reduces quadratically as we increase $r$ (the size of T-DAB’s time window) to values greater than $1$. Likewise, segmenting the $L$ input time steps to patches of a smaller length (in V-DAB) results in quadratic computational benefits.
Figure 3 depicts the GPU VRAM memory consumption of DeformTime during training compared to other transformer-based models w.r.t. the length of the look-back window ($L$; part A) and the number of input variables ($C+1$; part B). In these experiments, we set both the batch size and the hidden dimension $d$ to $64$. While assessing the impact of $L$, we set $C\!+\!1$ to $32$, and while assessing the impact of the amount of input variables, we set $L$ to $56$. We then present the average memory consumption across $50$ training epochs.
Figure 3: GPU memory (VRAM) consumption based on (A) the length (time steps) of the look-back window ($L$), and (B) the number of input variables ($C\!+\!1$).
Overall, DeformTime has a relatively small (GPU VRAM) memory footprint. As the length of the look-back window increases (from $28$ to $140$ time steps), iTransformer, TimeXer, and DeformTime are showing a stable memory consumption that does not exceed 111\!1 GB. We note that the low memory consumption of iTransformer / TimeXer is expected as they encode all input / exogenous input over the sequence dimension, which guarantees consistent GPU memory usage regardless of the value of $L$. In DeformTime, the input partitioning and adaptive transformations within the V-DAB and T-DAB modules reduce computational complexity, leading to reduced memory consumption, especially when $L$ is large. At the same time, the memory footprints of PatchTST and Crossformer linearly increase with $L$.
As we now increase the number of input variables from ($32$ to $512$), other methods display a linear increase in memory consumption (delayed and considerably lower for iTransformer and TimeXer compared to PatchTST and Crossformer), whereas for DeformTime memory consumption is almost unaffected and stable (e.g. $<\!1\!$ GB for DeformTime, but $>\!24\!$ GB for PatchTST). Hence, we argue that our method can handle a greater amount of exogenous predictors more efficiently.
Compared to other transformer-based MTS forecasting methods (Crossformer, PatchTST, and iTransformer), our method exhibits consistently low memory consumption as the number of input variables increases (vs. a linear increase for the other methods). Furthermore, increasing the look-back window does not overly affect memory consumption either (vs. a more accelerated linear increase for Crossformer and PatchTST). Hence, we deduce that DeformTime is adept to MTS forecasting tasks with a considerable amount of exogenous predictors and longer look-back windows.